# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 4. Essential and exotic groupoid Banach algebras
The reader interested only in Hausdorff groupoids may skip this section. When the groupoid $\mathcal{G}$ is non-Hausdorff we need to extend the previous definition of groupoid Banach algebras to allow constructions that ‘disregard’ the points of discontinuity for quasi-continuous functions in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$. In the $C^{*}$-algebraic context this quotient was introduced in [KM21] and examples go back (at least) to the work of Nekrashevych [Nek04], [Nek09]. The crucial observation is that elements of $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ are continuous modulo meager sets (i.e. countable unions of nowhere dense sets).
@@
For every $f\in\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ there is a meager set $M\subseteq\mathcal{G}$ such that $f$ is continuous at every point $\gamma\in\mathcal{G}\setminus M$.
@@
$\mathfrak{M}_{0}(\mathcal{G},\mathcal{L}):=\{f\in\mathfrak{C}_{0}(\mathcal{G},\mathcal{L}):\text{$\operatorname{supp}(f)$ is meager}\}$
where by support we mean the set
@@
We show that the elements in $\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$ have to be supported on a subgroupoid of $\mathcal{G}$ where Hausdorffness fails. To this end we adopt the nomenclature of Dixmier [Dix51].
@@
We apply the above definition to the groupoid $\mathcal{G}$. In particular, $\gamma\in\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ if there is $\eta\in\mathcal{G}\setminus\{\gamma\}$ such that $\gamma$ and $\eta$ cannot be separated by disjoint open sets (equivalently, there is a net $(\gamma_{i})_{i}\subseteq\mathcal{G}$ converging to $\gamma$ and $\eta$). The authors of [BM25] call arrows in $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ dangerous, which was inspired by [KM21, Definition 7.14] where the points in $X\setminus\mathcal{G}_{H}$ were given this name.
Statements similar to those in following lemma can also be found in [BM25], which appeared on arXiv six months after the first preprint of this article, and were proved independetely.
@@
Assume that $(\gamma_{i})_{i}\subseteq\mathcal{G}$ converges towards two different elements $\gamma,\gamma^{\prime}\in\mathcal{G}$. Then $(\gamma_{i}^{-1})_{i}\subseteq\mathcal{G}$ converges towards two different elements $\gamma^{-1},\gamma^{\prime-1}\in\mathcal{G}$, so $(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})^{-1}=\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$. Also $r(\gamma)=r(\gamma^{\prime})$ and $d(\gamma)=d(\gamma^{\prime})$, because $r,d$ are continuous and $X$ is Hausdorff. This in particular implies (1), cf. [KM21, Lemma 7.15]. Thus if $\eta\in\mathcal{G}$ is composable with $\gamma$ it is also composable with $\gamma^{\prime}$ and we may construct a net $(\eta_{i})_{i}$ convergent to $\eta$ such that $\eta_{i}$ and $\gamma_{i}$ are composable, so the net $(\eta_{i}\gamma_{i})_{i}$ converges to two different elements $\eta\gamma$ and $\eta\gamma^{\prime}$. Indeed, fixing two bisections $U,V\in\operatorname{Bis}(\mathcal{G})$ where $\eta\in V$, $\gamma\in U$ and $d(V)=r(U)$ we may assume that $(\gamma_{i})_{i}\subseteq U$, and then the elements $\eta_{i}:=d|_{V}^{-1}(r(\gamma_{i}))$ form the desired net. This shows that $\mathcal{G}(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})\subseteq\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$. By taking inverses, and using $(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})^{-1}=(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})$, we also get $(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})\mathcal{G}\subseteq(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})$. Since arrows in the groupoid are invertible, these relations for $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ are equivalent to saying that $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ and $\mathcal{G}_{\operatorname{H}}$ are full subgroupoids of $\mathcal{G}$. If $\mathcal{G}$ has a countable open cover consisting of bisections, then since open bisections are Baire spaces, $\mathcal{G}$ is a Baire space. Also by [KM21, Lemma 7.15], $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ is meager, and therefore it has empty interior. This proves item (2). Item (3) follows from (the proof of) [KM21, Lemma 7.13]. ∎
@@
A groupoid $\mathcal{G}$ is called principal if there are no points $x\in X$ with non-trivial isotropy, and $\mathcal{G}$ is called topologically principal if the set of points with non-trivial isotropy has empty interior. By Lemma 4.4(1), if $\mathcal{G}$ is principal then $\mathcal{G}=\mathcal{G}_{\operatorname{H}}$ is Hausdorff. If $\mathcal{G}$ is topologically principal then $\mathcal{G}_{\operatorname{H}}$ is dense in $\mathcal{G}$.
@@
    $\operatorname{supp}(f)$ is meager in $\mathcal{G}$, i.e. $f\in\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$;
@@
Implication (1$\Rightarrow$(2) holds because we can cover $\operatorname{supp}(f)$ with a countable family of bisections (cf. the proof of Lemma 4.1), and so its union is a Baire space. Implication (2$\Rightarrow$(3) is obvious. And (3) implies (1) because if $\operatorname{supp}(f)$ is not meager in $\mathcal{G}$, then by Lemma 4.1 there is a point $\gamma_{0}\in\operatorname{supp}(f)$ at which $f$ is continuous. Thus for $\varepsilon=|f(\gamma_{0})|/2$ the set $\{\gamma\in\mathcal{G}:|f(\gamma)|>\varepsilon\}$ contains a neighbourhood of $\gamma_{0}$ and so has non-empty interior. Hence (1)(3) are equivalent.
Implications (4$\Rightarrow$(5$\Rightarrow$(6) are clear. Since $(f^{*}*f)(x)=\sum_{\gamma\in d^{-1}(x)}|f(\gamma)|^{2}$ for $x\in X$, (6) implies (3). We close the cycle of implications by showing that (1) implies (6). To this end, note that $\operatorname{supp}(f)$ can be covered by a countable family $\{U_{n}\}_{n=1}^{\infty}\subseteq\operatorname{Bis}(\mathcal{G})$, because $f$ is a uniform limit of a sequence of functions in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$, which are supported on a finite union of bisections. So $\operatorname{supp}(f^{*}*f)\cap X=\bigcup_{n=1}^{\infty}d(\operatorname{supp}(f)\cap U_{n})$ is meager if each of the sets $\operatorname{supp}(f)\cap U_{n}$ is meager. This proves that (1)(6) are equivalent.
Lemma 4.4(3) implies that if $\operatorname{supp}(f)\cap\mathcal{G}_{\operatorname{H}}\neq\emptyset$ then $\operatorname{supp}(f)$ has non-empty interior. Thus (2) implies (7). The converse implication is trivial if we assume that $\mathcal{G}_{\operatorname{H}}$ is dense (which is guaranteed if there is a countable cover of $\mathcal{G}$ by bisections, by Lemma 4.4(2), or when $\mathcal{G}$ is topologically principal, see Remark 4.5). ∎
@@
In some pathological situations, see [KM21, Example 7.16], it may happen that $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ and so in general condition (7) above does not imply the others.
@@
If $\mathcal{G}$ is ample and every compact open set in $\mathcal{G}$ is regular open, i.e. it is equal to the interior of its closure, then $\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})=\{0\}$.
@@
Let $K(\mathcal{G},\mathcal{L})\subseteq\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ be the linear span of sections $g_{V}\in C(V,\mathcal{L})$ where $V$ is a compact open bisection and $|g_{V}|=const$ on $V$. Since $\mathcal{G}$ is ample, $K(\mathcal{G},\mathcal{L})$ is dense in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$. Indeed, the linear span of sections in $C_{0}(V,\mathcal{L})$, $V\in S(\mathcal{L})$, is dense in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$, and for every $V\in S(\mathcal{L})$ we have $C_{0}(V,\mathcal{L})\cong C_{0}(V)$, and it is well known that the linear span of indicator functions of compact open sets is dense in $C_{0}(V)$ when $V$ is totally disconnected.
We claim that for any $g\in K(\mathcal{G},\mathcal{L})$ and $c>0$ the set $|g|^{-1}(c)=\{\gamma\in\mathcal{G}:|g(\gamma)|=c\}$ is either empty or has non-empty interior. Indeed, we have $g=\sum_{V\in F}g_{V}$, where $F$ is a finite collection of compact open bisections and $g_{V}\in C(V,\mathcal{L})$ with $|g_{V}|=const$ on $V\in F$. By the inclusion-exclusion principle, putting for any non-empty $F_{0}\subseteq F$
@@
we get a family $\{V_{F_{0}}^{F}\}_{\emptyset\neq F_{0}\subseteq F}$ of pairwise-disjoint sets such that $V=\bigsqcup_{V\in F_{0}\subseteq F}V_{F_{0}}^{F}$, for every $V\in F$ (cf. [CFST14, Remark 2.5] or [BKM23, Lemma 6.2]). Thus $|g|$ is constant on each of the sets $V_{F_{0}}^{F}$. Hence if $|g|^{-1}(c)\neq\emptyset$, then $|g|^{-1}(c)$ contains a non-empty set $V_{F_{0}}^{F}$. But the set $V_{F_{0}}^{F}$ is the difference of two compact open sets $\bigcap_{V\in F_{0}}V$ and $\bigcup_{W\in F\setminus F_{0}}W$, which are regular open by assumption. Hence $V_{F_{0}}^{F}$ has a non-empty interior by [CEPSS19, Lemma 2.1].
Now let $f\in\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ be non-zero. Find $g\in K(\mathcal{G},\mathcal{L})$ with $\|f-g\|_{\infty}<\|f\|_{\infty}/3$, and so $\|g\|_{\infty}>2\|f\|_{\infty}/3$. Thus there is $\gamma\in\mathcal{G}$ with $c:=|g(\gamma)|>2\|f\|_{\infty}/3$. By the above claim $|g|^{-1}(c)$ has a non-empty interior. Moreover, $|g|^{-1}(c)\subseteq\operatorname{supp}(f)$ because for each $\gamma\in|g|^{-1}(c)$ we have $|f(\gamma)|\geq c-\|f\|_{\infty}/3>\|f\|_{\infty}/3>0$. Hence $\operatorname{supp}(f)$ has nonempty interior and therefore $f\notin\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$. ∎
@@ 4.1. Essential and exotic groupoid Banach algebras
For any open $U\subseteq\mathcal{G}$ we denote by $\mathfrak{M}(U,\mathcal{L})\subseteq\mathcal{B}(U,\mathcal{L})$ the space of bounded Borel sections of meager support, and we consider
@@
as the quotinet Banach space. We denote by $q_{U}:\mathcal{B}(U,\mathcal{L})\to\mathcal{D}(U,\mathcal{L})$ the quotient map. If $U,V\in\operatorname{Bis}(\mathcal{G})$ then the convolution and involution induce mappings $\mathcal{D}(U,\mathcal{L})\times\mathcal{D}(V,\mathcal{L})\to\mathcal{D}(UV,\mathcal{L})$ and $\mathcal{D}(U,\mathcal{L})\to\mathcal{D}(U^{*},\mathcal{L})$ that we will still denote by * and . Also we may and we will treat $\mathcal{D}(U,\mathcal{L})$ as a subspace of $\mathcal{D}(\mathcal{G},\mathcal{L})$. When the twist is trivial we will omit writing it. In particular,
@@
where the minimum exists because countable intersections of comeager sets are comeager. The Banach $C_{0}(X)$-bimodule structure on $\mathcal{B}(U,\mathcal{L})$ factors to a Banach $C_{0}(X)$-bimodule structure on $\mathcal{D}(U,\mathcal{L})$, and we have a $C_{0}(X)$-bimodule map $\operatorname{\mathbb{E}}_{U}:=q_{U}\circ E_{U}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{D}(U,\mathcal{L})$. Now we proceed as in Section 3, but with $E_{U}$ replaced by $\operatorname{\mathbb{E}}_{U}$. In the $C^{*}$-algebraic context $\operatorname{\mathbb{E}}_{X}$ was called a local expectation in [KM21].
@@
An exotic groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$ is a Hausdorff completion $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that $C_{0}(X)$ embeds isometrically into $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and $\operatorname{\mathbb{E}}_{X}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{D}(X)$ factors through a contractive map $\operatorname{\mathbb{E}}_{X}^{\mathcal{R}}:F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})\to\mathcal{D}(X)$, that is the corresponding seminorm satisfies $\|q_{X}(f|_{X})\|\leq\|f\|_{\mathcal{R}}$ for $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and $\|f\|_{\infty}=\|f\|_{\mathcal{R}}$ for $f\in C_{c}(X)$. If in addition $\operatorname{\mathbb{E}}_{X}^{\mathcal{R}}$ is faithful, we call $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ an essential groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$.
@@
A $C^{*}$-algebra is an exotic groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$ if and only if it is an exotic $C^{*}$-algebra of $(\mathcal{G},\mathcal{L})$ in the sense of [KM21]. There is a unique $C^{*}$-algebra which is an essential groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$, and this is the algebra introduced in [KM21, Definition 7.12] (for $\mathbb{F}=\mathbb{C}$). Every groupoid Banach algebra is an exotic groupoid Banach algebra and the converse holds when $\mathcal{G}$ is Hausdorff because then for every $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})=C_{c}(\mathcal{G},\mathcal{L})$ we have $f|_{X}\in C_{c}(X)$ and therefore $\|q_{X}(f|_{X})\|=\|f|_{X}\|_{\infty}$. In particular if $\mathcal{G}$ is Hausdorff, then the classes of essential and reduced groupoid Banach algebras coincide.
@@
The $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule structure on $\mathcal{B}(\mathcal{G},\mathcal{L})$ induced by convolution, defined in Lemma 3.4, factors to a $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule structure on $\mathcal{D}(\mathcal{G},\mathcal{L})$, which is continuous in the sense that for any $\{f_{n}\}_{n=1}^{\infty}\subseteq\mathcal{B}(\mathcal{G},\mathcal{L}),f\in\mathcal{B}(\mathcal{G},\mathcal{L})$ and $g\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ if $q_{\mathcal{G}}(f_{n})\to q_{\mathcal{G}}(f)$, then $q_{\mathcal{G}}(g*f_{n})\to q_{\mathcal{G}}(g*f)$ and $q_{\mathcal{G}}(f_{n}*g)\to q_{\mathcal{G}}(f*g)$. This makes $q_{\mathcal{G}}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{D}(\mathcal{G},\mathcal{L})$ a $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule map.
@@
We adapt the proof of Proposition 3.6. Implications (3$\Rightarrow$(2$\Rightarrow$(1) are clear (by putting $\operatorname{\mathbb{E}}^{\mathcal{R}}_{U}(f):=j_{\mathcal{R}}^{\operatorname{e}}(f)|_{U}$ for $f\in F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$, $U\in\operatorname{Bis}(\mathcal{G})$). We assume (1). For any $U\in S(\mathcal{L})$ and $f\in\mathfrak{C}_{c}(r(U)\mathcal{G},\mathcal{L})$, taking $b$ as in Lemma 3.5 we get

### v3

@@ 4. Essential and exotic groupoid Banach algebras
The reader interested only in Hausdorff groupoids may skip this section. When the groupoid $\mathcal{G}$ is non-Hausdorff we need to extend the previous definition of groupoid Banach algebras to allow constructions that ‘disregard’ the points of discontinuity for quasi-continuous functions in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$. In the $C^{*}$-algebraic context, this quotient was introduced in [KM21] and examples go back (at least) to the work of Nekrashevych [Nek04], [Nek09]. The crucial observation is that elements of $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ are continuous modulo meager sets (that is, countable unions of nowhere dense sets).
@@
For every $f\in\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$, there is a meager set $M\subseteq\mathcal{G}$ such that $f$ is continuous at every point $\gamma\in\mathcal{G}\setminus M$.
@@
4.1. Quasi-continous functions with meager support
@@
$\mathfrak{M}_{0}(\mathcal{G},\mathcal{L}):=\{f\in\mathfrak{C}_{0}(\mathcal{G},\mathcal{L}):\text{$\operatorname{supp}(f)$ is meager}\},$
where by $\operatorname{supp}$ we mean the strict support, given by
@@
We show that the elements in $\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$ have to be supported on a subgroupoid of $\mathcal{G}$ where Hausdorffness fails. To this end, we adopt the nomenclature of Dixmier [Dix51].
@@
We apply the above definition to the groupoid $\mathcal{G}$. In particular, $\gamma\in\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ if there is $\eta\in\mathcal{G}\setminus\{\gamma\}$ such that $\gamma$ and $\eta$ cannot be separated by disjoint open sets (equivalently, there is a net $(\gamma_{i})_{i}$ in $\mathcal{G}$ converging to $\gamma$ and $\eta$). The authors of [BM25] call arrows in $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ dangerous, which was inspired by [KM21, Definition 7.14] where the points in $X\setminus\mathcal{G}_{H}$ were given this name.
Statements similar to those in the following lemma can also be found in [BM25], which appeared on arXiv after the first preprint of this article, and were proved independently.
@@
Assume that a net $(\gamma_{i})_{i}$ in $\mathcal{G}$ converges towards two different elements $\gamma,\gamma^{\prime}\in\mathcal{G}$. Then $(\gamma_{i}^{-1})_{i}$ converges towards two different elements $\gamma^{-1},\gamma^{\prime-1}\in\mathcal{G}$, so $(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})^{-1}=\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$. Also $r(\gamma)=r(\gamma^{\prime})$ and $d(\gamma)=d(\gamma^{\prime})$, because $r,d$ are continuous and $X$ is Hausdorff. This in particular implies (1), cf. [KM21, Lemma 7.15]. Thus, if $\eta\in\mathcal{G}$ is composable with $\gamma$, it is also composable with $\gamma^{\prime}$ and we may construct a net $(\eta_{i})_{i}$ convergent to $\eta$ such that $\eta_{i}$ and $\gamma_{i}$ are composable, so the net $(\eta_{i}\gamma_{i})_{i}$ converges to two different elements $\eta\gamma$ and $\eta\gamma^{\prime}$. Indeed, fixing two bisections $U,V\in\operatorname{Bis}(\mathcal{G})$ where $\eta\in V$, $\gamma\in U$ and $d(V)=r(U)$, we may assume that $(\gamma_{i})_{i}$ is a net in $U$, and then the elements $\eta_{i}:=d|_{V}^{-1}(r(\gamma_{i}))$ form the desired net. This shows that $\mathcal{G}(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})\subseteq\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$. By taking inverses, and using $(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})^{-1}=(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})$, we also get $(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})\mathcal{G}\subseteq(\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}})$. Since arrows in the groupoid are invertible, these relations for $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ are equivalent to saying that $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ and $\mathcal{G}_{\operatorname{H}}$ are full subgroupoids of $\mathcal{G}$. If $\mathcal{G}$ has a countable open cover consisting of bisections, then since open bisections are Baire spaces, $\mathcal{G}$ is a Baire space. Also by [KM21, Lemma 7.15], $\mathcal{G}\setminus\mathcal{G}_{\operatorname{H}}$ is meager, and therefore it has empty interior. This proves item (2). Item (3) follows from (the proof of) [KM21, Lemma 7.13]. ∎
@@
Recall that a groupoid $\mathcal{G}$ is called principal if there are no points $x\in X$ with non-trivial isotropy, and $\mathcal{G}$ is called topologically principal if the set of points with non-trivial isotropy has empty interior. By Lemma 4.4(1), if $\mathcal{G}$ is principal then $\mathcal{G}=\mathcal{G}_{\operatorname{H}}$ is Hausdorff. If $\mathcal{G}$ is topologically principal, then $\mathcal{G}_{\operatorname{H}}$ is dense in $\mathcal{G}$.
@@
    $\operatorname{supp}(f)$ is meager in $\mathcal{G}$, that is, $f\in\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$;
@@
[…]
Lemma 4.4(3) implies that if $\operatorname{supp}(f)\cap\mathcal{G}_{\operatorname{H}}\neq\emptyset$, then $\operatorname{supp}(f)$ has non-empty interior. Thus (2) implies (7). The converse implication is trivial if we assume that $\mathcal{G}_{\operatorname{H}}$ is dense (which is guaranteed if there is a countable cover of $\mathcal{G}$ by bisections, by Lemma 4.4(2), or when $\mathcal{G}$ is topologically principal, see Remark 4.5). ∎
@@
In some pathological situations, see [KM21, Example 7.16], it may happen that $\mathcal{G}_{\operatorname{H}}=\emptyset$ and so in general condition (7) above does not imply the others.
@@
If $\mathcal{G}$ is ample and every compact open set in $\mathcal{G}$ is regular open, that is, it is equal to the interior of its closure, then $\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})=\{0\}$.
@@
Let $K(\mathcal{G},\mathcal{L})\subseteq\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ be the linear span of sections $g_{V}\in C(V,\mathcal{L})$ where $V$ is a compact open bisection and $|g_{V}|=\textit{const}$ on $V$. Since $\mathcal{G}$ is ample, $K(\mathcal{G},\mathcal{L})$ is dense in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$. Indeed, the linear span of sections in $C_{0}(V,\mathcal{L})$, $V\in S(\mathcal{L})$, is dense in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$, and for every $V\in S(\mathcal{L})$ we have $C_{0}(V,\mathcal{L})\cong C_{0}(V)$, and it is well known that the linear span of indicator functions of compact open sets is dense in $C_{0}(V)$ when $V$ is totally disconnected.
We claim that for any $g\in K(\mathcal{G},\mathcal{L})$ and $c>0$ the set $|g|^{-1}(c)=\{\gamma\in\mathcal{G}:|g(\gamma)|=c\}$ is either empty or has non-empty interior. Indeed, we have $g=\sum_{V\in F}g_{V}$, where $F$ is a finite collection of compact open bisections and $g_{V}\in C(V,\mathcal{L})$ with $|g_{V}|=\textit{const}$ on $V\in F$. By the inclusion-exclusion principle, putting for any non-empty $F_{0}\subseteq F$
@@
we get a family $\{V_{F_{0}}^{F}\}_{\emptyset\neq F_{0}\subseteq F}$ of pairwise-disjoint sets such that $V=\bigsqcup_{V\in F_{0}\subseteq F}V_{F_{0}}^{F}$, for every $V\in F$ (cf. [CFST14, Remark 2.5] or [BKM25, Lemma 6.2]). Thus $|g|$ is constant on each of the sets $V_{F_{0}}^{F}$. Hence if $|g|^{-1}(c)\neq\emptyset$, then $|g|^{-1}(c)$ contains a non-empty set $V_{F_{0}}^{F}$. But the set $V_{F_{0}}^{F}$ is the difference of two compact open sets $\bigcap_{V\in F_{0}}V$ and $\bigcup_{W\in F\setminus F_{0}}W$, which are regular open by assumption. Hence $V_{F_{0}}^{F}$ has a non-empty interior by [CEPSS19, Lemma 2.1].
Now let $f\in\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ be non-zero. Find $g\in K(\mathcal{G},\mathcal{L})$ with $\|f-g\|_{\infty}<\|f\|_{\infty}/3$, and so $\|g\|_{\infty}>2\|f\|_{\infty}/3$. Thus there is $\gamma\in\mathcal{G}$ with $c:=|g(\gamma)|>2\|f\|_{\infty}/3$. By the above claim $|g|^{-1}(c)$ has a non-empty interior. Moreover, $|g|^{-1}(c)\subseteq\operatorname{supp}(f)$ because, for each $\gamma\in|g|^{-1}(c)$, we have $|f(\gamma)|\geq c-\|f\|_{\infty}/3>\|f\|_{\infty}/3>0$. Hence $\operatorname{supp}(f)$ has nonempty interior and therefore $f\notin\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$. ∎
@@ 4.2. Essential and exotic groupoid Banach algebras
For any open $U\subseteq\mathcal{G}$, we denote by $\mathfrak{M}(U,\mathcal{L})\subseteq\mathcal{B}(U,\mathcal{L})$ the space of bounded Borel sections of meager support, and we consider
@@
as the quotient Banach space. Let $q_{U}:\mathcal{B}(U,\mathcal{L})\to\mathcal{D}(U,\mathcal{L})$ be the quotient map. If $U,V\in\operatorname{Bis}(\mathcal{G})$, then the convolution and involution induce mappings $\mathcal{D}(U,\mathcal{L})\times\mathcal{D}(V,\mathcal{L})\to\mathcal{D}(UV,\mathcal{L})$ and $\mathcal{D}(U,\mathcal{L})\to\mathcal{D}(U^{*},\mathcal{L})$ that we will still denote by * and . Also we may and we will treat $\mathcal{D}(U,\mathcal{L})$ as a subspace of $\mathcal{D}(\mathcal{G},\mathcal{L})$. When the twist is trivial we will omit writing it. In particular,
@@
where the minimum exists because countable intersections of comeager sets are comeager. The Banach $C_{0}(X)$-bimodule structure on $\mathcal{B}(U,\mathcal{L})$ factors to a Banach $C_{0}(X)$-bimodule structure on $\mathcal{D}(U,\mathcal{L})$, and we have a $C_{0}(X)$-bimodule map $\operatorname{\mathbb{E}}_{U}:=q_{U}\circ E_{U}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{D}(U,\mathcal{L})$. Now we proceed as in Section 3, but with $E_{U}$ replaced by $\operatorname{\mathbb{E}}_{U}$. In the $C^{*}$-algebraic context, $\operatorname{\mathbb{E}}_{X}$ was called a local expectation in [KM21].
@@
An exotic groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$ is a Hausdorff completion $F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})$ of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that $C_{0}(X)$ embeds isometrically into $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and $\operatorname{\mathbb{E}}_{X}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{D}(X)$ induces a contractive map $\operatorname{\mathbb{E}}_{X}^{\mathcal{R}}:F_{\mathcal{R}}({\mathcal{G},\mathcal{L}})\to\mathcal{D}(X)$, that is, the corresponding seminorm satisfies $\|q_{X}(f|_{X})\|\leq\|f\|_{\mathcal{R}}$ for $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and $\|f\|_{\infty}=\|f\|_{\mathcal{R}}$ for $f\in C_{c}(X)$. If in addition $\operatorname{\mathbb{E}}_{X}^{\mathcal{R}}$ is faithful, we call $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ an essential groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$.
@@
A $C^{*}$-algebra is an exotic groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$ if and only if it is an exotic $C^{*}$-algebra of $(\mathcal{G},\mathcal{L})$ in the sense of [KM21]. There is a unique $C^{*}$-algebra which is an essential groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$, and this is the algebra introduced in [KM21, Definition 7.12] (for $\mathbb{F}=\mathbb{C}$). Every groupoid Banach algebra is an exotic groupoid Banach algebra and the converse holds when $\mathcal{G}$ is Hausdorff because then, for every $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})=C_{c}(\mathcal{G},\mathcal{L})$, we have $f|_{X}\in C_{c}(X)$ and therefore $\|q_{X}(f|_{X})\|=\|f|_{X}\|_{\infty}$. In particular, if $\mathcal{G}$ is Hausdorff then the classes of essential and reduced groupoid Banach algebras coincide.
@@
The $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule structure on $\mathcal{B}(\mathcal{G},\mathcal{L})$ induced by convolution, defined in Lemma 3.4, factors to a $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule structure on $\mathcal{D}(\mathcal{G},\mathcal{L})$, which is continuous in the sense that for any sequence $(f_{n})_{n=1}^{\infty}$ in $\mathcal{B}(\mathcal{G},\mathcal{L})$, $f\in\mathcal{B}(\mathcal{G},\mathcal{L})$ and $g\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, if $q_{\mathcal{G}}(f_{n})\to q_{\mathcal{G}}(f)$, then $q_{\mathcal{G}}(g*f_{n})\to q_{\mathcal{G}}(g*f)$ and $q_{\mathcal{G}}(f_{n}*g)\to q_{\mathcal{G}}(f*g)$. This turns $q_{\mathcal{G}}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{D}(\mathcal{G},\mathcal{L})$ into a $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule map.
@@
[…]